OpenStudy (anonymous):

5(2/5 - 3/10) the 2/5 and 3/10 are fractions, btw.

OpenStudy (anonymous):

5(2/5-3/10) =5((4-3)/10) =5/10 =1/2

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

yw.